Rewriting the equation V = 2t + 1 t − 1 with t in terms of V , gives [1] t = 6 V . [2] t = V +1 V +2 . [3] t = 1−V V −2 . [4] t = V +1 V −2 . 


1
Expert's answer
2021-05-25T08:44:02-0400

Given equation is,


"V=\dfrac{2t+1}{t-1}"

Cross multiplying


"V(t-1)=2t+1\\\Rightarrow Vt-V=2t+1\\\Rightarrow Vt-2t=V+1"

Taking common factor 't' outside,


"\Rightarrow t(V-2)=V+1\\\Rightarrow t=\dfrac{V+1}{V-2}"

So, right answer is option [4] "\boxed{t=\dfrac{V+1}{V-2}}"